# Impact of the carbon price on credit portfolio’s loss with stochastic collateral

### v1

@@ 3.2 When collateral is a financial asset
Let note the collateral cash flows $(\overline{F}_{t}^{n})_{t\in\mathbb{R}_{+}}$, its dynamics is similar to the firm cash flows introduced in Assumption 2.5. We have for all $t\in\mathbb{R}_{+}$,
@@
and by (2.20), the approached collateral value as
@@
Therefore, the following proposition and its proof are inspired by Lemma 2.6 and corollary 2.7, gives a proxy of the collateral value.
@@ 3.3 When collateral is commercial or residential property
    the energy price $\mathfrak{f}(\cdot,\mathfrak{p})$ is non decreasing on $\mathbb{R}_{+}$ for all $\mathfrak{p}$.
@@
with $(\overline{\mathcal{W}}_{t})_{t\in\mathbb{R}_{+}}$ is a standard Brownian motion independent to $B^{\mathcal{Z}}$ introduced in Standing Assumption 2.1 and driving the productivity of the economy. Moreover, $C^{n}_{0}$, $r$, $R_{n},\overline{\sigma}>0$, $\rho\in\mathbb{R}_{+}^{I}$, and $\chi\in\mathcal{C}^{1}(\mathbb{R}_{+},\mathbb{R}_{+})$. We introduce the following filtration $\mathbb{U}:=(\mathcal{U}_{t})_{t\in\mathbb{R}^{*}}$ with for $t\geq 0$, $\mathcal{U}_{t}:=\sigma\left(\left\{\overline{W}_{s},B^{\mathcal{Z}}_{s}:s\leq t\right\}\right)$.
@@
with $\mathfrak{f}^{\mathfrak{p}}_{1},\mathfrak{f}^{\mathfrak{p}}_{0}>0$ and $\delta$ is the carbon price defined in the Standing Assumption 2.3 or an example given in (2.12).
@@
with $c_{0}>0$ and $c_{1}\geq-1$. This choice of $\mathfrak{c}$ allows us to model that when a building has a bad energy efficiency, its renovation is costly.
@@
By using the housing price under the climate transition as given in Proposition 3.9, we can then derive a precised expression of $\mathrm{LGD}$ when the collateral exists and is a building. We have:
@@ 3.4 Expected and Unexpected losses
    Given that $\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}=1-\gamma$ for $1\leq n\leq N_{1}$, to compute $\sum_{n=1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}\cdot\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$, all you have to do is calculate $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$.
@@
    Given that $N_{1}+1\leq n\leq N_{2}$, the collaterals are properties, therefore, to compute $\sum_{n=N_{2}+1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}^{n}_{t,T,\delta}\cdot\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$, we compute first $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$. Then we compute $\mathrm{LGD}^{n}_{t,T,\delta}$ directly through (3.37).
@@ 3.5 Remarks on the determinants of LGD
We could also look at the sensitivities of the LGD to each of these variables and parameters. However, the expressions of LGD we obtained are not very tractable so that it would be difficult to get detailed expressions of theses sensitivities. If necessary, they can be calculated using numerical methods.
@@ 4.1 Calibration and estimation
We will calibrate the model parameters on a set of data ranging from time $\mathfrak{t}_{0}$ to $\mathfrak{t}_{1}$. In practice, $\mathfrak{t}_{0}=1978$ and $\mathfrak{t}_{1}=t_{\circ}=2021$. From now on, we will discretize the observation interval into $M\in\mathbb{N}^{*}$ steps $t_{m}=\mathfrak{t}_{0}+\frac{\mathfrak{t}_{1}-\mathfrak{t}_{0}}{M}m$ for $0\leq m\leq M$. We note $\mathcal{T}^{M}:=\{t_{0},t_{1},\ldots,t_{M}\}$. We will not be interested in convergence results here.
@@ 4.1.2 Estimation of economic parameters
As in Galí, (2015), we assume a unitary Frisch elasticity of labor supply so $\varphi=1$ and the utility of consumption is logarithmic so $\sigma=1$, while we calibrate $({\boldsymbol{\lambda}}^{ij})_{i,j\in\mathcal{I}}$ and $({\chi}^{i})_{i\in\mathcal{I}}$in the same way as in Bouveret et al., (2023). We can then compute the functions $\chi$ and $\Lambda$ defined in Proposition 2.4, followed by the function $\widehat{v}^{i}$ as defined in (2.16). We can also compute the output growth $\left(\Delta^{{Y}}_{t_{m}}=(\log({Y}_{t_{m}}^{i})-\log({Y}_{t_{m-1}}^{i}))_{j\in\mathcal{I}}\right)_{1\leq m\leq M}$ directly from data.
Without carbon tax in any sector, it follows from (2.17) in Corollary 2.4 that, for each $1\leq m\leq M$, the computed consumption growth $\Delta^{{Y}}_{t_{m}}$ is equal to $\Delta^{{Y}}_{t_{m}}=\frac{\mathfrak{t}_{1}-\mathfrak{t}_{0}}{M}(\mathbf{I}_{I}-\widehat{\boldsymbol{\lambda}})^{-1}\widehat{\Theta}_{t_{m}}$ when $\mathbf{I}_{I}-\widehat{\boldsymbol{\lambda}}$ is not singular; hence $\widehat{\Theta}_{t_{m}}=\frac{M}{\mathfrak{t}_{1}-\mathfrak{t}_{0}}(\mathbf{I}_{I}-\widehat{\boldsymbol{\lambda}})\Delta^{{Y}}_{t_{m}}$. We can then compute the estimations $\widehat{\mu}$, $\widehat{\Gamma}$, $\widehat{\Sigma}$ and $\widehat{\varsigma}$, parameters $\mu$, $\Gamma$, $\Sigma$, and $\varsigma$ (all defined in Standing Assumption 2.1), as detailed in Sopgoui, (2024)[Section 3.1.1.].
@@ 4.1.3 Estimation of firm and of the credit model parameters
    Knowing the output growth $\left(\Delta^{{Y}}_{t}\right)_{t\in\mathcal{T}^{M}}$, we calibrate the factor loading $\mathfrak{a}_{n_{i}}$ and the standard deviation $\sigma_{n_{i}}$, according to Assumption 2.5, appealing to the regression
    $\sum_{n\in g_{i}}\log{F_{t_{m}}^{n}}-\log{F_{t_{m-1}}^{n}}=(\#g_{i})\mathfrak{a}^{n_{i}}\Delta^{{Y}}_{t_{k}}+\sqrt{\frac{\mathfrak{t}_{1}-\mathfrak{t}_{0}}{M}\#g_{i}}\sigma_{\mathfrak{b}^{n_{i}}}\mathfrak{u}_{t_{m}}\quad\text{where}\quad\mathfrak{u}_{t_{m}}\sim\mathcal{N}(0,1),\quad\forall\quad 1\leq m\leq M.$ (4.1)
@@
    We then estimate the barrier $B^{n_{i}}$ by MLE as detailed in (Gordy and Heitfield, , 2002, Section 3):
    we compute
@@
    with $D^{n_{i}}$ the Binomial random variable standing for the conditional number of defaults, and $\mathrm{PD}^{n_{i}}_{t_{m},1,0}$ in Proposition 2.29, depending on $\sigma_{\mathfrak{b}^{n_{i}}}=\widehat{\sigma}_{\mathfrak{b}^{n_{i}}}$, $\mathfrak{a}^{n_{i}}=\widehat{\mathfrak{a}}^{n_{i}}$, for $1\leq m\leq M$, $\delta_{t_{m}}=0$ and on $B^{n_{i}}$.
@@ 4.1.4 Calibration of collateral
    we calibrate the factor loading $\widehat{\overline{\mathfrak{a}}}_{n}$ and the standard deviation $\widehat{\overline{\sigma}}_{n}$, according to (3.12), appealing to the regression
    $\log{\overline{F}_{t_{m}}^{n}}-\log{\overline{F}_{t_{m-1}}^{n}}={\overline{\mathfrak{a}}}^{n}\Delta^{{Y}}_{t_{m}}+\sqrt{\frac{\mathfrak{t}_{1}-\mathfrak{t}_{0}}{M}}\overline{\sigma}_{n}\mathfrak{u}_{t_{m}}\quad\text{where}\quad\mathfrak{u}_{t_{m}}\sim\mathcal{N}(0,1),\quad\text{for all}\quad 1\leq m\leq M.$ (4.2)
@@ 4.2 Simulations
In this section as well, the idea here is not to (re)demonstrate or improve convergence results.
@@ 4.2.1 Of the productivities $\mathcal{Z}$ and $\mathcal{A}$
Let $K~{}\in\mathbb{N}$, for $0\leq k\leq K$, we note $u_{k}=t_{\circ}+\frac{t_{\star}-t_{\circ}}{K}k$ for $0\leq k\leq K$. We would like to simulate $\mathcal{Z}_{u_{k}}$ and $\mathcal{A}_{u_{k}}$. For $\mathcal{Z}$, we adopt the Euler-Maruyana Maruyama, (1955); Kanagawa, (1988) scheme as in Sopgoui, (2024)[Section 3.2.1].
@@ 4.2.2 Of the probability of over-indebtedness PD and of LGD
For $n\in\{1,\ldots,N\}$ and $t_{\circ}\leq t\leq t_{\star}$, We would like to compute $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$ as defined in (2.29) as well as $\mathrm{LGD}_{t,T,\mathfrak{d}}^{n}$ defined in (3.21) and $\mathrm{LGD}_{t,T,\delta}^{n}$ in (3.37). After simulating $\mathcal{Z}_{t}$ and $\mathcal{A}_{t}$ as described in 4.2.1, we get $\widehat{\mathcal{Z}}_{t}$ and $\widehat{\mathcal{A}}_{t}$. Then, for each $1\leq i\leq I$ and for each $n\in g_{i}$, we have
@@
    where $F^{n}_{0}$ and $\mathcal{D}_{t+T}^{n}$ are know, $\widehat{\mathfrak{d}}$ defined in Section 4.1.1, ^$\widehat{\mathfrak{R}}^{n}_{t+T}(\widehat{\mathfrak{d}}))$ in (2.22) in Theorem 2.6, $\widehat{\Gamma},\widehat{\varsigma},\widehat{v}$ in Section 4.1.2, $\widehat{\mathfrak{a}}^{n\cdot},\widehat{\sigma}^{n}$ in Section 4.1.3, $\widehat{\Upsilon}_{T}:=\widehat{\Gamma}^{-1}(\mathbf{I}_{I}-e^{-\widehat{\Gamma}T})$ and with $u_{l}:=\frac{Tl}{L},l=0,\ldots,L$.
@@ 4.2.3 Of the (un)expected losses EL and UL
$\begin{split}\widehat{\mathrm{EL}}^{N,T}_{t}&:=\sum_{n=1}^{N_{1}}(1-\gamma)\mathrm{EAD}_{t+T}^{n}\cdot\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}+\sum_{n=N_{1}+1}^{N_{2}}\mathrm{EAD}_{t+T}^{n}\cdot\widehat{\mathrm{LGD}}^{n}_{t,T,\mathfrak{d}}\cdot\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}\\ &\qquad\qquad+\sum_{n=N_{2}+1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\widehat{\mathrm{LGD}}^{n}_{t,T,\delta}\cdot\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}.\end{split}$ (4.8)
@@
$\widehat{\mathcal{Z}}_{t+T|t}^{p}|\mathcal{G}_{t}\sim\mathcal{N}\left(e^{-\widehat{\Gamma}T}\widehat{\mathcal{Z}}_{t},\frac{T}{L}\sum_{l=0}^{L}e^{-\widehat{\Gamma}u_{l}}\widehat{\Sigma}\widehat{\Sigma}^{\top}e^{-\widehat{\Gamma}^{\top}u_{l}}\right),$
@@
$\widehat{\mathcal{A}}_{t+T|t}^{p}|\mathcal{G}_{t}\sim\mathcal{N}\left(\widehat{\mu}T+\widehat{\varsigma}\widehat{\Upsilon}_{T}\widehat{\mathcal{Z}}_{t}+\widehat{\mathcal{A}}_{t},\widehat{\varsigma}^{2}\widehat{\Gamma}^{-1}\left[\frac{T}{L}\sum_{l=0}^{L}\left(e^{-\widehat{\Gamma}u_{l}}-\mathbf{I}_{I}\right)\widehat{\Sigma}\widehat{\Sigma}^{\top}\left(e^{-\widehat{\Gamma}u_{l}}-\mathbf{I}_{I}\right)\right](\widehat{\Gamma}^{-1})^{\top}\right),$

### v2

@@ 4.2 When collateral is a financial asset
Let note the collateral cash flows $(\overline{F}_{t}^{n})_{t\in\mathbb{R}_{+}}$, its dynamics is similar to the firm cash flows introduced in 3.6. We have for all $t\in\mathbb{R}_{+}$,
@@
and by (3.17), a proxy collateral value as
@@
Therefore, the following proposition (whose the proof is inspired by Lemma 3.7 and corollary 3.8) gives a proxy of the collateral value.
@@ 4.3 When collateral is commercial or residential property
    the energy price $\mathfrak{f}(\cdot,\mathfrak{p})$ is non decreasing on $\mathbb{R}_{+}$ for each type of energy $\mathfrak{p}$.
@@
with $(\overline{\mathcal{W}}_{t})_{t\in\mathbb{R}_{+}}$ is a standard Brownian motion independent to $B^{\mathcal{Z}}$ introduced in 3.1 and driving the productivity of the economy. Moreover, $C^{n}_{0}$, $r$, $R_{n},\overline{\sigma}>0$, $\rho\in\mathbb{R}_{+}^{I}$, and $\chi\in\mathfrak{D}(\mathbb{R}_{+},\mathbb{R}_{+})$. We introduce the following filtration $\mathbb{U}:=(\mathcal{U}_{t})_{t\in\mathbb{R}_{+}}$ with respect to $t\geq 0$, $\mathcal{U}_{t}:=\sigma\left(\left\{\overline{W}_{s},B^{\mathcal{Z}}_{s}:s\leq t\right\}\right)$.
@@
with $\mathfrak{f}^{\mathfrak{p}}_{1},\mathfrak{f}^{\mathfrak{p}}_{0}>0$ and $\delta$ is the carbon price defined in the 3.3 or an example given in (3.9).
@@
with $c_{0}>0$ and $c_{1}\geq-1$. This choice of $\mathfrak{c}$ allows us to model that when a building has a bad energy efficiency, its renovation is costly. This is similar to that of Sopgoui, (2024) and is inspired on the data given in Caumont et al., (2016), which summarizes the results of the survey on energy renovation of single-family homes in France.
@@
By using the housing price under the climate transition as given in theorem 4.9, we can then derive a precised expression of $\mathrm{LGD}$ when the collateral exists and is a building. We have:
@@ 4.4 Expected and Unexpected losses
    Given that $\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}=1-\gamma$ for $1\leq n\leq N_{1}$, to compute $\sum_{n=1}^{N_{1}}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}^{n}_{t,T,\mathfrak{d}}\cdot\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$, all you have to do is calculate $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$.
@@
    Given that $N_{2}+1\leq n\leq N$, the collaterals are properties, therefore, to compute $\sum_{n=N_{2}+1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\mathrm{LGD}^{n}_{t,T,\delta}\cdot\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$, we compute first $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$. Then we compute $\mathrm{LGD}^{n}_{t,T,\delta}$ directly through (4.37).
@@ 4.5 Remarks on the determinants of LGD
We could also look at the sensitivities of the LGD to each of these variables and parameters. However, the expressions of LGD we obtained are not very tractable so that it would be difficult to get detailed expressions of theses sensitivities. If necessary, they can be calculated using numerical methods (see Caprioli et al., (2023); Fermanian and Scaillet, (2005)).
@@ 5.1 Calibration and estimation
We will calibrate the model parameters on a set of data ranging from time $\mathfrak{t}_{0}$ to $\mathfrak{t}_{1}$. In practice, $\mathfrak{t}_{0}=1978$ and $\mathfrak{t}_{1}=t_{\circ}=2021$. From now on, we will discretize the observation interval into $M\in\mathbb{N}^{*}$ steps $t_{m}=\mathfrak{t}_{0}+\frac{\mathfrak{t}_{1}-\mathfrak{t}_{0}}{M}m$ for $0\leq m\leq M$. We note $\mathcal{T}^{M}:=\{t_{0},t_{1},\ldots,t_{M}\}$. We will not be interested in convergence results here. For the sake of clarity, we will omit the dependence of each estimated parameter/function on $\mathfrak{t}_{0}$ and $\mathfrak{t}_{1}$. We then note $\widehat{x}$ the estimation of $x$.
@@ 5.1.2 Estimation of economic parameters
As in Galí, (2015), we assume a unitary Frisch elasticity of labor supply so $\varphi=1$ and the utility of consumption is logarithmic so $\sigma=1$, while we calibrate $({\boldsymbol{\lambda}}^{ij})_{i,j\in\mathcal{I}}$ and  $({\psi}^{i})_{i\in\mathcal{I}}$ in the same way as in Bouveret et al., (2025). We can then compute the functions $\Psi$ and $\Lambda$ defined in proposition 3.5, followed by the function $\widehat{v}^{i}$ as defined in (3.13). We can also compute the output growth $\left(\Delta^{{Y}}_{t_{m}}=(\log({Y}_{t_{m}}^{i})-\log({Y}_{t_{m-1}}^{i}))_{j\in\mathcal{I}}\right)_{1\leq m\leq M}$ directly from data.
Without carbon tax in any sector, it follows from (3.14) in Corollary 3.5 that, for each $1\leq m\leq M$, the computed consumption growth $\Delta^{{Y}}_{t_{m}}$ is equal to $\Delta^{{Y}}_{t_{m}}=\frac{\mathfrak{t}_{1}-\mathfrak{t}_{0}}{M}(\mathbf{I}_{I}-\widehat{\boldsymbol{\lambda}})^{-1}\widehat{\Theta}_{t_{m}}$ when $\mathbf{I}_{I}-\widehat{\boldsymbol{\lambda}}$ is not singular; hence $\widehat{\Theta}_{t_{m}}=\frac{M}{\mathfrak{t}_{1}-\mathfrak{t}_{0}}(\mathbf{I}_{I}-\widehat{\boldsymbol{\lambda}})\Delta^{{Y}}_{t_{m}}$. We can then compute the estimations $\widehat{\mu}$, $\widehat{\Gamma}$, $\widehat{\Sigma}$, and $\widehat{\varsigma}$, of parameters $\mu$, $\Gamma$, $\Sigma$, and $\varsigma$ (all defined in 3.1), as detailed in Sopgoui, (2024)[Section 3.1.1.].
@@ 5.1.3 Estimation of firm and of the credit model parameters
    Knowing the output growth $\left(\Delta^{{Y}}_{t}\right)_{t\in\mathcal{T}^{M}}$, we calibrate the factor loading $\mathfrak{a}_{n_{i}}$ and the standard deviation $\sigma_{n_{i}}$, according to 3.6, appealing to the regression. For all $1\leq m\leq M$,
    $\sum_{n\in g_{i}}\log{F_{t_{m}}^{n}}-\log{F_{t_{m-1}}^{n}}=(\#g_{i})\mathfrak{a}^{n_{i}}\Delta^{{Y}}_{t_{k}}+\sqrt{\frac{\mathfrak{t}_{1}-\mathfrak{t}_{0}}{M}\#g_{i}}\sigma_{\mathfrak{b}^{n_{i}}}\mathfrak{u}_{t_{m}}\quad\text{where}\quad\mathfrak{u}_{t_{m}}\sim\mathcal{N}(0,1).$ (5.1)
@@
    We then obtain the barrier $B^{n_{i}}$ by maximum likelihood estimation (MLE) as detailed in (Gordy and Heitfield, , 2002, Section 3). We have
@@
    with $D^{n_{i}}$ the Binomial random variable standing for the conditional number of defaults, and $\mathrm{PD}^{n_{i}}_{t_{m},1,0}$ in eq. 3.25, depending on $\sigma_{\mathfrak{b}^{n_{i}}}=\widehat{\sigma}_{\mathfrak{b}^{n_{i}}}$, $\mathfrak{a}^{n_{i}}=\widehat{\mathfrak{a}}^{n_{i}}$, for $1\leq m\leq M$, $\delta_{t_{m}}=0$ and on $B^{n_{i}}$.
@@ 5.1.4 Calibration of collateral
    We calibrate the factor loading $\widehat{\overline{\mathfrak{a}}}_{n}$ and the standard deviation $\widehat{\overline{\sigma}}_{n}$, according to (4.12), appealing to the regression. For all $1\leq m\leq M$
    $\log{\overline{F}_{t_{m}}^{n}}-\log{\overline{F}_{t_{m-1}}^{n}}={\overline{\mathfrak{a}}}^{n}\Delta^{{Y}}_{t_{m}}+\sqrt{\frac{\mathfrak{t}_{1}-\mathfrak{t}_{0}}{M}}\overline{\sigma}_{n}\mathfrak{u}_{t_{m}}\quad\text{where}\quad\mathfrak{u}_{t_{m}}\sim\mathcal{N}(0,1).$ (5.2)
@@ 5.2 Simulations
In this section as well, the idea here is not to (re)demonstrate or improve convergence results. Recall the climate transition should start at $t_{\circ}$ and end at $t_{\star}$. Therefore, we would like to compute the risk measures for $t\in[t_{\circ},t_{\star}]$.
@@ 5.2.1 Of the productivities $\mathcal{Z}$ and $\mathcal{A}$
Let $K~{}\in\mathbb{N}^{*}$, for $0\leq k\leq K$, we note $u_{k}=t_{\circ}+\frac{t_{\star}-t_{\circ}}{K}k$ for $0\leq k\leq K$. We would like to simulate $\mathcal{Z}_{u_{k}}$ and $\mathcal{A}_{u_{k}}$. For $\mathcal{Z}$, we adopt the Euler-Maruyana (see Maruyama, (1955); Kanagawa, (1988)) scheme as in Sopgoui, (2024)[Section 3.2.1].
@@ 5.2.2 Of the probability of over-indebtedness PD and of LGD
For $n\in\{1,\ldots,N\}$ and $t_{\circ}\leq t\leq t_{\star}$, We would like to compute $\mathrm{PD}^{n}_{t,T,\mathfrak{d}}$ as defined in (3.25) as well as $\mathrm{LGD}_{t,T,\mathfrak{d}}^{n}$ defined in (4.21) and $\mathrm{LGD}_{t,T,\delta}^{n}$ in (4.37). After simulating $\mathcal{Z}_{t}$ and $\mathcal{A}_{t}$ as described in 5.2.1, we get $\widehat{\mathcal{Z}}_{t}$ and $\widehat{\mathcal{A}}_{t}$. Then, for each $1\leq i\leq I$ and for each $n\in g_{i}$, we have
@@
    where $F^{n}_{0}$ and $\mathcal{D}_{t+T}^{n}$ are know, $\widehat{\mathfrak{d}}$ defined in section 5.1.1, ^$\widehat{\mathfrak{R}}^{n}_{t+T}(\widehat{\mathfrak{d}}))$ in (3.19) in proposition 3.7, $\widehat{\Gamma},\widehat{\varsigma},\widehat{v}$ in section 5.1.2, $\widehat{\mathfrak{a}}^{n\cdot},\widehat{\sigma}^{n}$ in section 5.1.3, $\widehat{\Upsilon}_{T}:=\widehat{\Gamma}^{-1}(\mathbf{I}_{I}-e^{-\widehat{\Gamma}T})$ and with $u_{l}:=\frac{Tl}{L},l=0,\ldots,L$.
@@ 5.2.3 Of the (un)expected losses EL and UL
$\begin{split}\widehat{\mathrm{EL}}^{N,T}_{t,\mathfrak{d}}&:=\sum_{n=1}^{N_{1}}(1-\gamma)\mathrm{EAD}_{t+T}^{n}\cdot\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}+\sum_{n=N_{1}+1}^{N_{2}}\mathrm{EAD}_{t+T}^{n}\cdot\widehat{\mathrm{LGD}}^{n}_{t,T,\mathfrak{d}}\cdot\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}\\ &\qquad\qquad+\sum_{n=N_{2}+1}^{N}\mathrm{EAD}_{t+T}^{n}\cdot\widehat{\mathrm{LGD}}^{n}_{t,T,\delta}\cdot\widehat{\mathrm{PD}}^{n}_{t,T,\mathfrak{d}}.\end{split}$ (5.8)
@@
$\widehat{\mathcal{Z}}_{t+T|t}^{p}|\mathcal{G}_{t}\sim\mathcal{N}\left(e^{-\widehat{\Gamma}T}\widehat{\mathcal{Z}}_{t},\frac{T}{L}\sum_{l=0}^{L}e^{-\widehat{\Gamma}u_{l}}\widehat{\Sigma}\widehat{\Sigma}^{\top}e^{-\widehat{\Gamma}^{\top}u_{l}}\right),$ (5.10)
@@
$\widehat{\mathcal{A}}_{t+T|t}^{p}|\mathcal{G}_{t}\sim\mathcal{N}\left(\widehat{\mu}T+\widehat{\varsigma}\widehat{\Upsilon}_{T}\widehat{\mathcal{Z}}_{t}+\widehat{\mathcal{A}}_{t},\widehat{\varsigma}^{2}\widehat{\Gamma}^{-1}\left[\frac{T}{L}\sum_{l=0}^{L}\left(e^{-\widehat{\Gamma}u_{l}}-\mathbf{I}_{I}\right)\widehat{\Sigma}\widehat{\Sigma}^{\top}\left(e^{-\widehat{\Gamma}u_{l}}-\mathbf{I}_{I}\right)\right](\widehat{\Gamma}^{-1})^{\top}\right),$ (5.11)
@@
5.3 Summary of the process
More concretely, the goal is to calculate the expected and unexpected losses at each $t\in[\mathfrak{t}_{0},\mathfrak{t}_{1}]$. To achieve that, we use (1) the number of firms rated $r_{t}$ and defaulted $d_{t}$ in each $t\in\mathcal{T}^{M}$, (2) all the firms’ cash flows $(F_{t}^{n})_{1\leq n\leq N}$ in each $t\in\mathcal{T}^{M}$, (3) for each borrower $n$, if a collateral exists and is a financial asset, its cash flows $(\overline{F}_{t}^{n})_{t\in\mathcal{T}^{M}}$ and if it is a building, its surface $R_{n}$ and its price at $\mathfrak{t}_{1}$, $C_{0}^{n}$, (4) the macroeconomic variables as well as the carbon intensities by sector, and the real estate index observed $REI_{t}$ in each $t\in\mathcal{T}^{M}$, and (5) the carbon price dynamics $(\delta_{t})_{t\in\{t_{\circ},\ldots,t_{\star}\}}$ given by the regulator. We proceed as follows.
  1. 1.
    From the macroeconomic historical data, we estimate the productivity parameters $\widehat{\Gamma}$, $\widehat{\mu}$$\widehat{\Sigma}$ and $\widehat{\varsigma}$, as well as the elasticities $\widehat{\psi}$ and $\widehat{\boldsymbol{\lambda}}$ as described in section 5.1.2. We also estimate the real estate market parameters $\widehat{\varrho}$, $\widehat{\vartheta}$, $\widehat{\nu}$, $\widehat{\overline{\sigma}}$, and $\widehat{\rho}$ as described in section 5.1.4.
  2. 2.
    For each $i\in\{1,\ldots,I\}$, we estimate the parameters $B^{n_{i}}$, $\sigma_{\mathfrak{b}^{n_{i}}}$, $\mathbf{a}^{n_{i}}$ using section 5.1.3, yielding $\widehat{B}^{n_{i}}$, $\widehat{\sigma}_{\mathfrak{b}^{n_{i}}}$, $\widehat{\mathbf{a}}^{n_{i}}$.
  3. 3.
    We compute the carbon price dynamics $(\delta_{t})_{t_{\circ}\leq t\leq t_{\star}}$ and the carbon intensities $(\tau_{t})_{t_{\circ}\leq t\leq t_{\star}}$, $(\zeta_{t})_{t_{\circ}\leq t\leq t_{\star}}$, and $(\kappa_{t})_{t_{\circ}\leq t\leq t_{\star}}$ as defined in 3.3 as well as the emissions cost rate $(\mathfrak{d}_{t})_{t_{\circ}\leq t\leq t_{\star}}$ defined in (3.8) and the output carbon cost function $v$ defined in (3.13).
  4. 4.
    For $t_{\circ}\leq t\leq t_{\star}$, we fix a large enough integer $D$, and simulate $D$ paths of the productivity process $(\widehat{\mathcal{Z}}_{t}^{d})_{1\leq d\leq D}$, then we derive $((\mathcal{A}_{t}^{\circ})^{d})_{1\leq d\leq D}$ as defined in section 5.2.1. Then, for each $p$, we simulate $D$ paths $(\widehat{\mathcal{Z}}_{t+T|t}^{d^{\prime}}|\mathcal{G}_{t}^{d})_{1\leq d\leq D}$ and $(\widehat{\mathcal{Z}}_{t+T|t}^{d^{\prime}}|\mathcal{G}_{t}^{d})_{1\leq d^{\prime}\leq D}$ using (5.10) and (5.11) respectively.
  5. 5.
    For each $n\in\{1,\ldots,N\}$ and for each $1\leq d,d^{\prime}\leq D$,
    • (a)
      we compute the probability of default $\widehat{\mathrm{PD}}_{t+1,\mathfrak{d}}^{n}$ using (5.3).
    • (b)
      Depending on the nature of the collateral, we compute $\widehat{\mathrm{LGD}}_{t+1,\delta}^{n}$ , using either (4.11), (5.6), or (5.7)
  6. 6.
    We finally compute the expected (resp. unexpected) losses $\widehat{\mathrm{EL}}^{N}_{t,\mathfrak{d}}$ (resp. $\widehat{\mathrm{UL}}^{N,T}_{t,\delta,\alpha}$), for each $t_{\circ}\leq t\leq t_{\star}$, using (5.8) (resp. (5.12)).